# The Madelung Problem of Finite Crystals

### v1

@@ Abstract
The classical Madelung problem concerns the electrostatic potential at the site of a reference ion in an infinite ionic crystal (e.g., NaCl or CsCl)Madelung (1918). When scaled by the potential of an isolated nearest-neighbor counterion, this defines the Madelung constant—a dimensionless quantity central to the lattice energy per ion pair. Its computation, however, is mathematically subtle: the conditionally convergent nature of the Coulomb series renders naive direct summation (e.g., over spherical or cubic shells) not only slow but also ambiguous, as the result may depend on the summation orderde Leeuw et al. (1980); Smith (1981).
@@
Table 1: Madelung constants for CsCl (${\mathbf{v}}_{1}=(0.5,0.5,0.5)$) computed using the explicitly corrected direct sum [Eqs. (1) to (3)] and the CS method [Eqs. (4) and (5)], for varying supercell sizes [^2] The reference value is 1.76 267 477 307 098csc .
@@
In this work, we overcome the slow convergence of conventional direct summation by incorporating higher-order multipolar effects via exact boundary and finite-size corrections—without resorting to charge or distance renormalization. For the CsCl-like two-particle configuration depicted in Fig. 1, the direct Coulomb sum for a finite crystal comprises of a shape- and size-independent bulk term $\nu_{\rm pbc}({\mathbf{r}})$, a size-independent boundary contribution $\nu_{\rm b}({\mathbf{r}}|{\mathbf{s}})$, and a finite-size correction $\nu_{\rm corr}({\mathbf{r}},p|{\mathbf{s}})$
@@
where ${\mathbf{r}}=(x,y,z)$ denotes the relative displacement, ${\mathbf{n}}=(n_{x},n_{y},n_{z})$ is the lattice vector, $n=|{\mathbf{n}}|$, and ${\cal L}(p|{\mathbf{s}})$ the set of all lattice vectors in a finite crystal of linear extent $p$ (with $K=2p+1$ unit cells per dimension) and shape $\mathbf{s}$}. The precise meaning of $\mathbf{s}$} will be clarified later. In the present explicit finite-size correction (EC) method, the Madelung constant is obtained by isolating the bulk contribution $\nu_{\rm pbc}$ from the direct sum $\nu({\mathbf{r}},p|{\mathbf{s}})$ and scaling it by the potential of an isolated counterion at the nearest-neighbor distance.
@@
Table 2 shows that, the correction scheme maintains high accuracy (better than $10^{-8}$ at $p=20$) across a range of displacement vectors, confirming its robustness for arbitrary displacements. As detailed later, the two-ion configuration serves as a fundamental building block for the Madelung problem: for crystals with multi-atom unit cells (e.g., the crystal structures in Tab. 3), the total Madelung constant is obtained by linearly superposing contributions from all symmetry-inequivalent ion pairs, each evaluated via the direct sum in Eq. (1) and corrected analytically using $\nu_{\rm b}+\nu_{\rm corr}$. This additive property, combined with the ${\cal O}(p^{-4})$ convergence of every term, enables a truly analytical, renormalization-free evaluation of Madelung constants—even for complex ionic structures such as NaCl, ZnS, CaF2 and CaTiO3. In the following, we provide rigorous derivations of $\nu_{\rm b}({\mathbf{r}}|{\mathbf{s}})$ and $\nu_{\rm corr}({\mathbf{r}},p|{\mathbf{s}})$ for a general orthogonal lattice. While $\nu_{\rm b}({\mathbf{r}}|{\mathbf{s}})$ admits a closed-form expression for crystals with arbitrary aspect ratios, the explicit closed-form expression $\nu_{\rm corr}({\mathbf{r}},p|{\mathbf{s}})$ in Eq. (3) is derived for cubic crystals in a cubic lattice. Nevertheless, this formulation is sufficient to compute highly accurate Madelung constants for a broad set of benchmark structures—requiring only minimal computational effort.
@@
(0, 0, 0)(0.5, 0, 0.5) (0.5, 0, 0)(0, 0.5, 0)   (0.25, 0.25, 0.25)(0.75, 0.25, 0.75)  (0.25, 0.25, 0.75)(0.25, 0.75, 0.25)
(0.5, 0.5, 0)(0, 0.5, 0.5)   (0, 0, 0.5)(0.5, 0.5, 0.5)   (0.75, 0.75, 0.25)(0.25, 0.75, 0.75)  (0.75, 0.25, 0.25)(0.75, 0.75, 0.75)
@@
This expression—and analogous sums for the electrostatic energy—constitutes the starting point for deriving the Ewald summation and the associated shape-dependent boundary contributions in the limit of a macroscopic (infinite) crystal (e.g.de Leeuw et al. (1980); Smith (1981); Ballenegger (2014); Hu (2014)). In contrast, we rigorously define the geometry of finite crystals for a general orthogonal lattice, thereby enabling a transparent and systematic treatment of finite-size effects. In the absence of such a formulation, the boundary contribution in the infinite limit remains ambiguous—often obscured by implicit assumptions and lacking direct justification from finite, physically realizable systems.
@@
allowing the potential to be recast in a pairwise formHu (2014); Yi et al. (2017); Zhao and Hu (2025)
@@
In prior work, the infinite-system limit of this interaction is decomposed asHu (2014); Yi et al. (2017); Zhao and Hu (2025)
@@
where $\nu_{\rm pbc}({\mathbf{r}})$ is geometry-independent and represents the bulk contribution under periodic boundary conditions (PBC). This term has served as a foundation for analytical studies of structural correlations in bulk electrolytes and dielectricsHu (2022); Gao et al. (2023), as well as dielectric responses at interfacesPan et al. (2017, 2019). The second term, $\nu_{\rm b}({\mathbf{r}}|{\mathbf{s}})$ accounts for the arrangement of charges at the infinite boundaryZhao and Hu (2025). Crucially, as we demonstrate below, it remains a non-negligible contribution even for finite crystals. In reciprocal space, it can be expressed asHu (2014); Zhao and Hu (2025)
@@
where ${\rm atan}(\cdot)$ is the inverse tangent function and the dimensionless parameter $\xi_{\alpha}$ are defined by normalizing each side length with the geometric mean of the three:
@@
For a cubic crystal with unit lattice constant ($l_{x}=l_{y}=l_{z}=1$), $\nu_{\rm b}({\mathbf{r}}|{\mathbf{s}})$ of Eq. (17) simplifies to Eq. (2).
@@
$\nu_{\rm corr}({\mathbf{r}},p)=-\sum_{{\mathbf{n}}\notin{\cal L}(p|{\mathbf{s}})}^{{\cal L}(\infty|{\mathbf{s}})}\left[\frac{1}{\left|{\mathbf{r}}+{\mathbf{n}}\right|}-\frac{1}{n}\right],$ (19)
@@
where $\cos\theta={\mathbf{r}}\cdot{\mathbf{n}}/(nr)$, $P_{k}(\cos\theta)$ is the $k$-th Legendre polynomial, and $\nabla_{\mathbf{n}}$ is the gradient operator with respect to the lattice vector $\mathbf{n}$}. Gelle and Lepetit performed a similar multipole expansion using spherical coordinates rather than Cartesian coordinatesGellé and Lepetit (2008). Their approach facilitates asymptotic estimation of truncation errors and analysis of convergence behavior. In contrast, we derive exact leading-order terms of $\nu_{\rm corr}({\mathbf{r}},p|{\mathbf{s}})$ directly in Cartesian coordinates.
@@
and analogously for $n_{y}^{2}$ and $n_{z}^{2}$. As a result, both the continuum approximation and the residual discrete dipolar terms vanish. Interestingly, Eq. (22) also holds for spherical truncation of cubic unit cells. Hence, the leading correction in both cubic and spherical crystals with $l_{x}=l_{y}=l_{z}$ stems solely from the $k=4$ multipole. In the SI, we evaluate the corresponding continuum integral explicitly for cubic crystals, obtaining the closed-form expression given in Eq. (3). Such a closed-form expression for the spherical case, however, remains unknown.

### v2

# arXiv:2512.18138v2[cond-mat.mtrl-sci] 28 Feb 2026

@@ Abstract
I Introduction
The classical Madelung problem concerns the electrostatic potential at the site of a reference ion in an infinite ionic crystal (e.g., NaCl or CsCl)Madelung (1918). When scaled by the potential of an isolated nearest-neighbor counterion, this defines the Madelung constant—a dimensionless quantity central to the lattice energy per ion pair. Its computation, however, is mathematically subtle: the conditionally convergent nature of the Coulomb series renders naive direct summation (e.g., over spherical or cubic shells) not only slow but also ambiguous, as the result may depend on the summation order (e.g.de Leeuw et al. (1980); Smith (1981, 2008); Ballenegger (2014); Hu (2014a); Pan et al. (2017); Zhao and Hu (2025)).
@@ I Introduction
Table 1: Madelung constants for CsCl (${\mathbf{v}}_{1}=(0.5,0.5,0.5)$) computed using the explicitly corrected direct sum [Eqs. (1) to (3)] and the CS method [Eqs. (4) and (5)], for varying supercell sizes. The reference value is 1.76 267 477 307 098csc .
@@
In this work, we resolve the long-standing ambiguity in the Madelung problem or the Coulomb lattice sum—stemming from the conditional convergence of the electrostatic lattice sum—by rigorously defining and analytically separating boundary and finite-size effects in finite crystals. Our approach achieves rapid convergence from conventional direct summation by incorporating higher-order multipolar effects via exact boundary and finite-size corrections—without resorting to charge or distance renormalization. For the CsCl-like two-particle configuration depicted in Fig. 1, the direct Coulomb sum for a finite crystal comprises of a shape- and size-independent bulk term $\nu_{\rm pbc}({\mathbf{r}})$, a size-independent boundary contribution $\nu_{\rm b}({\mathbf{r}}|{\mathbf{s}})$, and a finite-size correction $\nu_{\rm corr}({\mathbf{r}},p|{\mathbf{s}})$
@@
where ${\mathbf{r}}=(x,y,z)$ denotes the relative displacement, ${\mathbf{n}}=(n_{x},n_{y},n_{z})$ is the lattice vector, $n=|{\mathbf{n}}|$, and ${\cal L}(p|{\mathbf{s}})$ the set of all lattice vectors in a finite crystal of linear extent $p$ (with $K=2p+1$ unit cells per dimension) and shape $\mathbf{s}$}. Here ${\mathbf{s}}=(0,0,0)$ indicates the shape of the finite cubic crystal. The meaning of $\mathbf{s}$} in a general orthogonal lattice will be clarified later. In the present explicit finite-size correction (EC) method, we isolate $\nu_{\rm pbc}$, which is the bulk contribution under the periodic boundary condition (PBC), from the direct sum $\nu({\mathbf{r}},p|{\mathbf{s}})$ and then scale it by the potential of an isolated counterion at the nearest-neighbor distance to obtain the Madelung constant. The EC method thus provides a well-defined approximation to the size- and shape-independent bulk term $\nu_{\rm pbc}$ by evaluating a finite Coulomb lattice sum supplemented with explicit boundary and finite-size corrections. Eq. (1) captures the essential physical meaning of the Madelung constant as originally conceived by EwaldEwald (1921), namely the electrostatic potential at a reference ion due to all other ions in the crystal—distinct from the response of the crystal to an externally applied fieldRayleigh (1892). It has long been recognized that conditionally convergent lattice sums can yield different results depending on the summation order or the geometric perspective adoptedde Leeuw et al. (1980). The key advance of the present work lies in its unambiguous definition of both size and shape for a finite crystal. In contrast, previous approaches—such as those based on infinite lattice sums (e.g., Ref.Hu (2014a) and references therein)—leave the boundary condition ill-defined, thereby obscuring the physical origin of shape-dependent contributions.
@@
Table 2 shows that, the correction scheme maintains high accuracy (better than $10^{-8}$ at $p=20$) across a range of displacement vectors, confirming its robustness for arbitrary displacements. As detailed later, the two-ion configuration serves as a fundamental building block for the Madelung problem: for crystals with multi-atom unit cells (e.g., the crystal structures in Tab. 3), the total Madelung constant is obtained by linearly superposing contributions from all symmetry-inequivalent ion pairs, each evaluated via the direct sum in Eq. (1) and corrected analytically using $\nu_{\rm b}+\nu_{\rm corr}$. This additive property, combined with the ${\cal O}(p^{-4})$ convergence of every term, enables a truly analytical, renormalization-free evaluation of Madelung constants—even for complex ionic structures such as NaCl, ZnS, CaF2 and CaTiO3.
@@
(0, 0, 0) (0.5, 0, 0) (0.25, 0.25, 0.25) (0.25, 0.25, 0.75)
(0.5, 0, 0.5) (0, 0.5, 0) (0.75, 0.25, 0.75) (0.25, 0.75, 0.25)
(0.5, 0.5, 0) (0, 0, 0.5) (0.75, 0.75, 0.25) (0.75, 0.25, 0.25)
(0, 0.5, 0.5) (0.5, 0.5, 0.5) (0.25, 0.75, 0.75) (0.75, 0.75, 0.75)
In the following, we first illustrate the essential meaning of Eq. (1) for a simple cubic lattice by comparing two cubic crystals in Section II and by comparing cubic and cuboid crystals in Section III, respectively. Rigorous derivations of $\nu_{\rm b}({\mathbf{r}}|{\mathbf{s}})$ and $\nu_{\rm corr}({\mathbf{r}},p|{\mathbf{s}})$ for a general orthogonal lattice are provided in Section IV. While $\nu_{\rm b}({\mathbf{r}}|{\mathbf{s}})$ admits a closed-form expression for crystals with arbitrary aspect ratios, the explicit closed-form expression $\nu_{\rm corr}({\mathbf{r}},p|{\mathbf{s}})$ in Eq. (3) is derived for cubic crystals in a cubic lattice. Nevertheless, Section V demonstrates that this formulation is sufficient to compute highly accurate Madelung constants for a broad set of benchmark structures—requiring only minimal computational effort. Finally, we draw conclusions in Section VI and provide evaluations of selected integrals in the Appendix and in the Supporting Information.
II Difference between two cubic crystals
Figure 2: Cross-sectional view of two cubic crystals of size $p$, each composed of unit charges ($\pm 1$), separated by the vectors $(t,0,0)$ (top) and (t1,0,0)(t-1,0,0) (middle), respectively. In the limit of large $p$, the difference of charges in their outermost layers can be approximated by a pair of uniformly charged parallel plates located at the planes x=(p+1t)x=-(p+1-t) and $x=p+t$ (bottom).
Eq. (1) decomposes the lattice sum over long-range Coulomb interactions into three distinct components: (i) the bulk term $\nu_{\rm pbc}({\mathbf{r}})$ which is a periodic function of $\mathbf{r}$}, (ii) the boundary term $\nu_{\rm b}({\mathbf{r}}|{\mathbf{s}})$, which depends on the shape of the crystal and is not a periodic function of $\mathbf{r}$}, (iii) the finite-size correction $\nu_{\rm corr}({\mathbf{r}},p|{\mathbf{s}})$, which vanishes at the limit $p\to\infty$. Obviously, if the interactions were short-ranged—decaying sufficiently fast or vanishing beyond a certain distance—the boundary term would not appear. In such cases, the bulk term accounts for the influence of nearby particles, while the finite-size correction captures the effect of the finite extent of the crystal. The shape-dependent boundary term has attracted considerable attention in the literature (see, e.g.de Leeuw et al. (1980); Smith (1981, 2008); Ballenegger (2014); Hu (2014a)).Previous derivations often relied on intricate integral transforms. In contrast, we present here a transparent derivation of the boundary effect using only elementary calculus and basic electrostatics, based on a direct comparison of two finite crystals.
To this end, we first consider two cubic crystals in a simple cubic lattice with unit lattice constant. Their relative displacements are taken as ${\mathbf{r}}_{1}=(t,0,0)$ and its periodic translation: ${\mathbf{r}}_{2}=(t-1,0,0)$, respectively. Both crystals have identical dimensions $(2p+1)\times(2p+1)\times(2p+1)$, differing only in the outermost layers perpendicular to the $x$ direction, as illustrated in Fig. 2. For large $p$, the finite-size correction becomes negligible, and the local environment around the central ion (marked by dot in Fig. 2) is identical in both crystals except at the boundaries. When the central ion is placed at the origin, the difference in the electrostatic potential it experiences arises solely from the two layers located at $x=p+t$ and x=(p+1t)x=-(p+1-t). Because $p$ is large, these layers can be approximated as uniformly charged square plates of side length $a=2p+1$ and charge densities $\sigma=\pm 1$.
The electrostatic potential produced by a single square plate of size $a$ and charge density $\sigma$ at a point on its central axis at distance $z$ is given by $\sigma\,\phi(z,a)$ with Hummer (1996); Ciftja (2020)
$\phi(z,a)=2a\,\psi(2z/a),$ (6)
where the dimensionless function $\psi(x)$ is defined as
$\psi(x)=\log\frac{\sqrt{x^{2}+2}+1}{\sqrt{x^{2}+2}-1}-x\,{\rm atan}\frac{1}{x\sqrt{x^{2}+2}},$ (7)
where ${\rm atan}(\cdot)$ is the inverse tangent function. With $z_{1}=p+1-t$ and $z_{2}=p+t$ for $t\in(0,1)$, the boundary contribution can be expressed as
$\nu_{\rm b}({\mathbf{r}}_{2}|{\mathbf{s}})-\nu_{\rm b}({\mathbf{r}}_{1}|{\mathbf{s}})=\lim_{p\to\infty}\phi(z_{1},a)-\phi(z_{2},a)$ (8)
To verify this relation, note that
$\frac{2z_{1}}{a}=1+\frac{1-2t}{a};\quad\frac{2z_{2}}{a}=1-\frac{1-2t}{a}.$ (9)
Expanding $\psi(x-\tau)$ around $x=0$ to second order yields
$\psi(x+\tau)=\psi(\tau)-x\,{\rm atan}\frac{1}{\tau\sqrt{\tau^{2}+2}}\\ +\frac{x^{2}}{\left(\tau^{2}+1\right)\sqrt{\tau^{2}+2}}+\cdots.$ (10)
Substituting $\tau=1$ and $x=\pm(1-2t)/a$, we obtain
$\lim_{p\to\infty}\phi(z_{1},a)-\phi(z_{2},a)=-4a\frac{\pi}{6}\frac{1-2t}{a}=\frac{2\pi}{3}\left(2t-1\right).$ (11)
This result matches the difference obtained directly from the boundary term in Eq. (2):
$\nu_{\rm b}({\mathbf{r}}_{2}|{\mathbf{s}})-\nu_{\rm b}({\mathbf{r}}_{1}|{\mathbf{s}})=-\frac{2\pi}{3}\left[(t-1)^{2}-t^{2}\right]=\frac{2\pi}{3}\left(2t-1\right).$ (12)
Thus, the boundary effect derived from the electrostatics of parallel plates agrees exactly with the analytical expression, confirming the validity of the approach.
III Difference between cubic and cuboid crystals
Figure 3: Cross-sectional view of cubic (top) and cuboid (middle) crystals of size $p$, both composed of unit charges ($\pm 1$), separated by the vectors $(t,0,0)$. For large $p$, the additional layers presented in the cuboid crystal can be approximated by pairs of uniformly charged parallel plates located at the planes x=(3p+1)x=-(3p+1), (3p+1t)-(3p+1-t), $\cdots$, (p+1)-(p+1), (p+1t)-(p+1-t), $p+1$, $p+1+t$, $\cdots$, $3p+1$, $3p+1+t$ (bottom).
By comparing two cubic crystals, the preceding section provided a tutorial demonstration that the boundary term $\nu_{\rm b}({\mathbf{r}}|{\mathbf{s}})$ is not a periodic function of $\mathbf{r}$}. In this section, we extend the analysis to compare cubic and cuboid crystals in a simple cubic lattice with unit lattice constant. This comparison illustrates that the boundary term also depends on the shape of the crystal.
Both crystals are displaced by the same vector ${\mathbf{r}}_{1}=(t,0,0)$. The cubic crystal of size $p$ is identical to the one considered previously. The cuboid crystal, by contrast, has dimensions $3(2p+1)\times(2p+1)\times(2p+1)$, i.e., it is three times longer in the $x$-direction. Fig. 3 shows a cross-sectional view of the two crystals in the $xy$-plane, along with their difference approximated as pairs of parallel plates.
For large $p$, each of the $(2p+1)$ additional layers in the cuboid crystal—indexed by $k=p+1,p+2,\cdots,3p+1$—can be represented by two pairs of oppositely charged parallel plates. Each plate is perpendicular to the $x$-axis, has side length $a=2p+1$, and carries a charge density $\sigma=\pm 1$. The contribution from the $k$-th layer consists of four terms: $2a\psi_{+}(k)$, $2a\psi_{-}(k)$, $2a\psi_{0}(k)$, $2a\psi_{0}(k)$ where
$\psi_{+}(k)=\psi\left(\frac{2k+2t}{a}\right),$ (13)
$\psi_{-}(k)=\psi\left(\frac{2k-2t}{a}\right),$ (14)
$\psi_{0}(k)=\psi\left(\frac{2k}{a}\right).$ (15)
and $\psi(x)$ is defined in Eq. (7).
Let ${\mathbf{s}}_{1}$ and ${\mathbf{s}}_{2}$ denote the shapes of the cubic and cuboic crystals, respectively. The boundary effect is fully captured by the following relation:
$\nu_{\rm b}({\mathbf{r}}_{1}|{\mathbf{s}}_{2})-\nu_{\rm b}({\mathbf{r}}_{1}|{\mathbf{s}}_{1})\\ =\lim_{p\to\infty}\sum_{k=p+1}^{3p+1}2a\left[\psi_{+}(k)+\psi_{-}(k)-2\psi_{0}(k)\right].$ (16)
Substituting $\tau=2k/a$ and $x=\pm 2t/a$ into the Taylor series in Eq. (10), the right hand side (rhs) of Eq. (16) becomes
$\text{rhs.}=\lim_{p\to\infty}\sum_{k=p+1}^{3p+1}\frac{16at^{2}}{\left(4k^{2}+a^{2}\right)\sqrt{4k^{2}+2a^{2}}}.$ (17)
Replacing the sum by an integral with the substitution $x=2k/a$ yields,
$\text{rhs.}=8t^{2}\int_{1}^{3}\frac{dx}{{\left(x^{2}+1\right)\sqrt{x^{2}+2}}}.$ (18)
As shown in the Appendix, the indefinite integral evaluates to
$\int\frac{dx}{{\left(x^{2}+1\right)\sqrt{x^{2}+2}}}={\rm atan}\frac{x}{\sqrt{x^{2}+2}}+\text{const.},$ (19)
so that
$\nu_{\rm b}({\mathbf{r}}_{1}|{\mathbf{s}}_{2})-\nu_{\rm b}({\mathbf{r}}_{1}|{\mathbf{s}}_{1})=8t^{2}\left[{\rm atan}\frac{3}{\sqrt{11}}-{\rm atan}\frac{1}{\sqrt{3}}\right].$ (20)
This result explicitly confirms that the boundary term depends on the crystal shape. The validity of the above equation will be further verified in the next section, where we derive the explicit expression for the boundary term of a cuboid crystal.
IV General results for an orthogonal lattice
@@
This expression—and analogous sums for the electrostatic energy—constitutes the starting point for deriving the Ewald summation and the associated shape-dependent boundary contributions in the limit of a macroscopic (infinite) crystal (e.g.de Leeuw et al. (1980); Smith (1981); Ballenegger (2014); Hu (2014a)). In contrast, we rigorously define the geometry of finite crystals for a general orthogonal lattice, thereby enabling a transparent and systematic treatment of finite-size effects. In the absence of such a formulation, the boundary contribution in the infinite limit remains ambiguous—often obscured by implicit assumptions and lacking direct justification from finite, physically realizable systems.
@@
allowing the potential to be recast in a pairwise formHu (2014a); Yi et al. (2017); Zhao and Hu (2025)
@@
In prior work, the infinite-system limit of this interaction is decomposed asHu (2014a); Yi et al. (2017); Zhao and Hu (2025)
@@
where $\nu_{\rm pbc}({\mathbf{r}})$ is geometry-independent and represents the bulk contribution. This term has served as a foundation for analytical studies of structural correlations in bulk electrolytes and dielectricsHu (2022); Gao et al. (2023), as well as dielectric responses at interfacesHu (2014b); Pan et al. (2017, 2019). It has been demonstrated recently thatZhao and Hu (2025), the pairwise sum over $\nu_{\rm pbc}({\mathbf{r}})$ provides critical insight into the role of the neutralizing background charge and enables a consistent formulation of energies and pressure in systems comprising both discrete and continuous charge distributionsOnegin et al. (2024); Li et al. (2025); Demyanov et al. (2025). The second term, $\nu_{\rm b}({\mathbf{r}}|{\mathbf{s}})$ accounts for the arrangement of charges at the infinite boundaryZhao and Hu (2025). Crucially, as we demonstrate below, it remains a non-negligible contribution even for finite crystals. In reciprocal space, it can be expressed asHu (2014a); Zhao and Hu (2025)
@@
where the dimensionless parameter $\xi_{\alpha}$ are defined by normalizing each side length with the geometric mean of the three:
@@
For a cubic crystal with unit lattice constant ($l_{x}=l_{y}=l_{z}=1$), $\nu_{\rm b}({\mathbf{r}}|{\mathbf{s}})$ of Eq. (32) simplifies to Eq. (2). Inserting Eq. (32) into the left hand side of Eq. (20) confirms the validity of Eq. (20) (see also Eq. (A9) in the Appendix).
@@
$\nu_{\rm corr}({\mathbf{r}},p|{\mathbf{s}})=-\sum_{{\mathbf{n}}\notin{\cal L}(p|{\mathbf{s}})}^{{\cal L}(\infty|{\mathbf{s}})}\left[\frac{1}{\left|{\mathbf{r}}+{\mathbf{n}}\right|}-\frac{1}{n}\right],$ (34)
@@
where $\cos\theta={\mathbf{r}}\cdot{\mathbf{n}}/(nr)$, $P_{k}(\cos\theta)$ is the $k$-th Legendre polynomial, and $\nabla_{\mathbf{n}}$ is the gradient operator with respect to the lattice vector $\mathbf{n}$}. While multipolar series underpins fast algorithms for evaluating potentials in Coulomb systemsGreengard and Rokhlin (1987); Greengard (1988); Kudin and Scuseria (1998), we instead provide an analytical treatment of the boundary term and asymptotic finite-size behavior for finite crystals with explicitly specified shape and size. Gelle and Lepetit performed a similar multipole expansion using spherical coordinates rather than Cartesian coordinatesGellé and Lepetit (2008). Their approach facilitates asymptotic estimation of truncation errors and analysis of convergence behavior. In contrast, we derive exact leading-order terms of $\nu_{\rm corr}({\mathbf{r}},p|{\mathbf{s}})$ directly in Cartesian coordinates.
@@
and analogously for $n_{y}^{2}$ and $n_{z}^{2}$. As a result, Eq. (36) completely vanishes, i.e., both the continuum approximation and the residual discrete dipolar terms vanish. Interestingly, Eq. (37) also holds for spherical truncation of cubic unit cells. Hence, the leading correction in both cubic and spherical crystals with $l_{x}=l_{y}=l_{z}$ stems solely from the $k=4$ multipole. In the SI, we evaluate the corresponding continuum integral explicitly for cubic crystals, obtaining the closed-form expression given in Eq. (3). Such a closed-form expression for the spherical case, however, remains unknown.
V Applications to various lattices